OpenStudy (anonymous):

if a bus travels at rate of 50 mph to get to new jersey from new york and a rate of 60 mph to get back to new york and the time taken was 32 hours, how long does it take to get back to new york

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

im trying!! let me try to google it

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

oh this is driving me crazy! It's in my head, I just can't seem to get it together!

OpenStudy (anonymous):

i know what you mean its driving me crazy too

OpenStudy (anonymous):

what do i do to get the answer though

OpenStudy (anonymous):

let x be the time he takes to go to new york. 32-x is the time he takes to go to new jersey

OpenStudy (anonymous):

I am confused that is 32hrs is that the time for round trip or just once?

OpenStudy (anonymous):

distance is the same

OpenStudy (anonymous):

i have the 32-x

OpenStudy (anonymous):

its round trip

OpenStudy (anonymous):

so speed x time must be the same for both

OpenStudy (anonymous):

x times 60 = (32-x) times 50

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

try reading my posts from the start and see if you understand

OpenStudy (anonymous):

i dnt understand can you write the steps together

OpenStudy (anonymous):

ok sure give me a moment

OpenStudy (anonymous):

let x be the time he takes to go back to new york so, time he takes to go to new jersey is (32-x) the distance is the same for both journeys. so, (speed x time) is the same x times 60 = (32-x) times 50

OpenStudy (anonymous):

\[60 \times x = (32-x) \times 50\]

OpenStudy (anonymous):

distance from new york to new jersey = distance from new jersey to new york speed to new jersey x time taken to new jersey = speed to new york x time taken to new york

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

the time to get back to NY is 14.55, so is it 55 minutes out of an hour or 55% of 60 min guys?

OpenStudy (anonymous):

where did u get 192 and 11 from (O.o) IM LOST

OpenStudy (anonymous):

it is 14.55 hrs

OpenStudy (anonymous):

irkiz how did u get that

OpenStudy (anonymous):

solve the equation i gave for x

OpenStudy (anonymous):

r*t=d Let t be the time taken to travel from NJ to NY. Solve the following for t:\[50(32-t)=60t\]\[t=\frac{160}{11}=14.55 \text{ hours}\]

OpenStudy (anonymous):

distance from new york to new jersey = distance from new jersey to new york speed to new jersey x time taken to new jersey = speed to new york x time taken to new york let time to go to NY be x time taken to go to NJ is (32-x) (32-x) times 50 = x times 60 solve for x

OpenStudy (anonymous):

OH I GET IT THANK YOI

OpenStudy (anonymous):

YOU*

OpenStudy (anonymous):

Ok i just had 2 paragraphs of explanation! Glad this question is OVER!!! Thank you irkiz!!

OpenStudy (anonymous):

you're welcome